### v2

# PAOLI: Pose-free Articulated Object Learning from Sparse-view Images

Abstract
We present a methodology to model articulated objects using a sparse set of images with unknown poses.
Current methods require dense multi-view observations and ground-truth camera poses. Our approach operates with as few as four views per articulation and no camera supervision. Our central insight is to first solve a robust correspondence and alignment problem between unaligned reconstructions, before part motions can be analyzed. We first reconstruct each articulation independently using recent advances in sparse-view 3D reconstruction, then learn a deformation field that establishes dense correspondences across poses. A progressive disentanglement strategy further separates static from moving parts, enabling robust separation of camera and object motion. Finally, we optimize geometry, appearance, and kinematics jointly with a self-supervised loss that enforces cross-view and cross-pose consistency. Experiments on the standard benchmark and real-world examples demonstrate that our method produces accurate and detailed articulated object representations under significantly weaker input assumptions than existing approaches.
1 Introduction
Automating human capabilities, such as reasoning about and interacting with pivoting doors, sliding drawers, opening boxes, remains an open problem. Practical solutions for modeling such articulated objects—rigid objects coupled via joints that allow rotation or translation— from video is beneficial to downstream applications such as robotic manipulation. Current approaches require dense, multi-view images across articulation poses, making them impractical. In this work, we show that articulated object modeling is possible with only four unposed images per articulation state, vastly reducing capture requirements.
Early approaches [42, 18, 48] rely on rich supervision such as 3D object shapes, articulation parameters, or part segmentations. While effective, this approach depends on high-quality 3D scans and labor-intensive manual annotations, which are both costly and difficult to scale. Recent methods reduce supervision by learning from unlabeled multiple views; rather than relying on manual labels, these approaches infer object structure and kinematics from observed motion [30, 10, 16]. For instance, PARIS [30] and ArticulateYourNeRF [10] follow an analysis-by-synthesis approach, learn object representations from pairs of multi-view images, using Neural Radiance Fields (NeRF) [38]. These models jointly recover 3D geometry and appearance while separating moving parts from static ones. ArticulatedGS [16] further improves efficiency and accuracy by replacing NeRF with Gaussian Splatting (GS) [21].
Despite these advances, the state of the art still relies on dense, paired observations, requiring around 100 paired views of the same object across two articulation states. Such dense capture is rarely practical. Moreover, these methods assume that all capture lie in a shared, pre-calibrated coordinate system, i.e., that the object’s static base is perfectly aligned across articulation states via known camera poses. In practice, this level of calibration is costly and rarely achievable, severely limiting deployment.
Figure 1: Given few unposed views of an articulated object across different articulated poses, PAOLI (our method) reconstructs its 3D geometry and appearance, while estimating part segmentation and joint axis of its moving parts (shown as arrows) in 3D. The reconstructed object can be rendered from novel viewpoints and articulation configurations.
We instead study articulated object reconstruction and modeling from unposed and sparse views, significantly more realistic yet fundamentally harder setting. Traditional 3D reconstruction methods such as structure-from-motion (SFM) [45, 46] or neural-based methods [53, 55] assume static scenes and cannot directly handle object motion. Treating each articulation state independently bypasses this issue but produces reconstructions in different coordinate frames, creating a severe misalignment problem. Articulated reasoning becomes impossible without first establishing a reliable 3D-to-3D correspondence between these unaligned shapes.
A common approach is to match sparse keypoints. However, keypoints are unreliable under large articulations, on textureless surfaces, and provide insufficient coverage for dense part segmentation. This makes them unsuitable for the experimental setting we target.
We introduce PAOLI, a methodology that learns dense correspondences via a optimized deformation field. Leveraging sparse view 3D reconstruction [62, 14], we first obtain independent GS reconstructions for each articulation state. A dense deformation field is optimized to warp the source model into alignment with the target state by enforcing 3D geometric consistency and 2D photometric appearance. Dense correspondences between the original source model and the target model are then established by using 1-Nearest-Neighbor (1-NN) to match the deformed source model to the target model. Finally to separate rigid parts, we introduce an iterative robust estimation algorithm that clusters inliers from the correspondence set using a pose estimator [59], thereby isolating static and moving components and recovering their kinematic parameters.
In summary, our main contributions are:
  1. 1.
    Problem formulation: We are the first to address articulated object reconstruction from sparse, unposed images, identifying the 3D-to-3D correspondence problem between unaligned reconstructions as the central challenge.
  2. 2.
    Dense correspondence via deformation: We propose a robust correspondence detector, using a dense deformation field, which is optimized with 3D and 2D losses. This approach outperforms keypoint-based correspondence methods and other adapted baselines.
  3. 3.
    A new customized solution: We demonstrate the first complete pipeline to successfully recover full articulated models including geometry, parts, and kinematics from sparse, and unposed data.
These contributions demonstrate that a robust solution to this correspondence problem is achieved not by sparse keypoints or pre-trained models, but by a dense deformation field optimized at test-time for each object, leveraging a combination of 3D geometry and 2D rendering losses.
2 Related Work
Pose-free 3D reconstruction
Classical Structure from Motion (SfM) [1, 34, 45, 47] jointly estimates camera poses and scene geometry via image registration, triangulation, and bundle adjustment. Recent learning-based multi-view stereo (MVS) methods [60, 55, 27, 53] achieve high-quality reconstructions without explicit feature matching. DUST3R [55] uses transformers to predict dense point maps and poses from image pairs, though bundle adjustment is still required. MASt3R [27] improves scalability with coarse-to-fine matching for high-resolution inputs, while VGGT [53] extends this paradigm with multi-task learning to jointly predict point maps, poses, depth, and tracks. These methods, however, focus on static scenes and do not explicitly model articulated motion or object parts.
Sparse View Gaussian Splats
Neural rendering [38, 21, 17, 54, 41] enables photorealistic novel view synthesis by rendering from learned 3D representations. More recently, 3D GS [21] has emerged as particularly compelling, offering efficient, high-quality representations [12, 8, 37, 25, 58]. Standard GS requires dense multi-view images with accurate poses. To relax this, some works use depth priors [29, 63, 9] or diffusion guidance [32, 22, 36], while others directly predict Gaussian parameters with feed-forward networks [61, 62, 49, 50, 4, 6]. Splatter Image [49] and LGM [50] handle single images, while MVSPlat [6] and PixelSplat [4] target sparse multi-view with known poses. FreeSplatter [62] and NoPoSplat [61] remove pose supervision by jointly predicting poses and Gaussians, but remain limited to static scenes without articulated motion. UFV-splatter [14] improves robustness to image viewpoints by finetuning FreeSplatter with images from unfavorable views, which enhances reconstruction and camera pose estimation.
Articulated Object Reconstruction
Reconstructing articulated objects is challenging due to coupled geometry and kinematics. Some methods focus on structure or motion alone [19, 23, 5, 13, 33, 20], while others use generative part-level graphs [31, 26]. More recent works jointly model geometry and appearance through optimization [40, 51, 56, 30, 10, 57, 35, 16]. DITTO [19] as one of the prioneer work that aims to model two-part articulated objects with point cloud input data and assumes aligned coordinate frame with pre-trained feed-forwrd model. PARIS [30] reconstructs parts and motions from dense multi-view scans with known poses; ArticulateYourNeRF [10] improves robustness via iterative segmentation and motion learning; while ArticulatedGS [16], ArtGS [35], and DTA [57] enhance quality with GS and depth priors. However, all rely on dense multi-view inputs and calibrated poses. Our approach instead reconstructs articulated objects from sparse, unposed views by leveraging recent feed-forward MVS advances.
3D-3D Correspondence and Matching
Finding 3D-to-3D correspondences is a fundamental problem. Approaches based on sparse keypoints, from classical descriptors like FPFH [44] to modern learned detectors [28, 64]. Recent registration methods for 3D Gaussian Splatting (GS) like GaussReg [3], proposed to establish coarse 3D keypoint matching between GS, which still struggles with accurate registration in 3D data, making them unsuitable for dense articulated motion analysis. A more robust alternative is to learn dense correspondences via deformation fields [15, 11, 2, 24, 7]. These approaches, however, typically rely on feed-forward, pre-trained models to predict deformation and are designed for meshes or point clouds, not 3D GS. In contrast, our method introduces a test-time, optimization-based framework to learn a dense deformation field specifically for GS, allowing us to solve the articulated motion problem. We then use a robust iterative algorithm to segment these dense correspondences into distinct rigid parts and their kinematics.
3 Method
We address the task of reconstructing articulated objects, defined as a composition of rigid parts connected by 1-DoF revolute or prismatic joints. Given two sets of images, each capturing an articulated object from sparse and unknown viewpoints (with a fixed pose within each set but different poses across sets), our objective is to reconstruct a ”digital twin” of the object. This model includes high-fidelity 3D geometry, appearance, and articulation parameters, enabling physically plausible dynamic rendering.
Our method consists of three main stages. First, we reconstruct the 3D geometry and estimate camera parameters for each image set independently. Next, we learn a deformation field to align the geometry and establish initial dense correspondences. Finally, we jointly optimize the geometry, part segmentation, and articulation parameters. An overview is shown in Fig. 2. Since the final optimization is intertwined and requires good initialization, the first two stages are crucial.
Figure 2: Overview of our method pipeline. Given two sets of K-view images (K=4) of an articulated object in different states, our approach reconstructs high-fidelity 3D geometry and estimates articulation parameters through a three-stage process: (1) initialization of Gaussian splats and camera parameters (top left), (2) Gaussian alignment via deformation fields (top right), and (3) joint optimization of geometry and articulation parameters (bottom).
3.1 Initialization of 3D Gaussian Splats
Given two sets of images, a source set $\mathcal{I}^{s}=\{I^{s}_{k}\}_{k=1}^{K}$ and a target set $\mathcal{I}^{t}=\{I^{t}_{k}\}_{k=1}^{K}$, each containing $K$ (e.g.4) pose-free views of the object, we first construct initial 3D Gaussian splats $G^{s}$ and $G^{t}$ independently. Each set $G=\{\mathbf{g}_{i}\}_{i=1}^{N}$ consists of $N$ Gaussian points, where each point $\mathbf{g}=(\bm{\mu},\mathbf{q},\mathbf{s},o,\mathbf{c})$ is parameterized by its mean $\bm{\mu}$, rotation $\mathbf{q}$, scale $\mathbf{s}$, opacity $o$, and color $\mathbf{c}$.
Since FreeSplatter is designed for static scenes, we apply it independently to the source set $\mathcal{I}^{s}$ and the target set $\mathcal{I}^{t}$. This process yields our two initial 3D Gaussian Splatting models, $G^{s}$ and $G^{t}$, each in its own arbitrary coordinate system. This step also provides initial camera parameters (both extrinsics and intrinsics) for each set. We then compute a single, averaged intrinsic matrix from the two sets and use this fixed intrinsic parameter for all subsequent optimization steps.
However, the initial Gaussians and extrinsic estimates are not perfectly pixel-aligned. To refine this, we freeze the GS and optimize only the extrinsic parameters ($\Pi^{s},\Pi^{t}$) for each set by minimizing a photometric rendering loss. This refinement provides a better initialization for the subsequent steps.
3.2 Gaussian Correspondences via Deformation Fields
A central challenge in building digital twins of articulated objects from pose-free images lies in aligning the canonical pose across different observation sets. As argued in our related work, a common approach for recovering pose is to match sparse keypoints [28, 64]. However, this is often unreliable for dense analysis and struggles with texture-less surfaces. A more robust alternative is to find a dense alignment. A naive strategy would be to optimize the mean and rotation of each Gaussian point independently, but this breaks local rigidity, since it allows unconstrained updates of individual parameters, whereas true object motion should remain rigid within each part.
To address this, we learn a per-scene deformation field, parameterized by a lightweight feed-forward network, $\mathcal{F}_{\text{deform}}$. This network is optimized from scratch for each object and predicts rigid transformations for Gaussian points in $G^{s}$ using their positional encodings, producing a deformed target representation $\hat{G}^{t}$. By constraining the network to generate low-frequency deformations, we enforce local rigidity while still capturing global motion. Specifically, $\mathcal{F}_{\text{deform}}$ maps the positional encoding of Gaussian means to a rigid transformation consisting of a rotation, translation, and scaling factor:
$\{\mathbf{R},\mathbf{t},s\}=\mathcal{F}_{\text{deform}}(\text{PE}(\bm{\mu}))$ (1)
where $\mathbf{R}\in\text{SO}(3)$ is parameterized by Euler angles $\in\mathbb{R}^{3}$ and converted to rotation matrices, $\mathbf{t}\in\mathbb{R}^{3}$ represents the translation vector, $s\in\mathbb{R}$ is a global scaling factor shared across all points, and $\text{PE}(\bm{\mu})$ is the positional encoding of the Gaussian means. This transformation is then applied to the source Gaussian points to obtain the deformed target representation.
The means of the source Gaussian points $\bm{\mu}\in\mathbb{R}^{3}$ are encoded as:
$\text{PE}(\bm{\mu})=[\sin(2^{k_{0}}\pi\bm{\mu}),\cos(2^{k_{0}}\pi\bm{\mu})]$ (2)
where $k_{0}$ represents the base frequency parameter. The parameter $k_{0}$ directly influences the rigidity of the deformation field: smaller values result in lower-frequency encodings that naturally constrain the deformation field to be more rigid, effectively guiding the network’s capacity to learn low-frequency non-rigid transformations.
Deformation Network Training
To optimize the parameters of $\mathcal{F}_{\text{deform}}$, we minimize the Chamfer distance $\mathcal{L}_{\text{CD}}$ between the means of the deformed source Gaussian $\hat{G}^{t}$ and the target one ${G}^{t}$, as well as $L1$ photometric loss $\mathcal{L}_{\text{photo}}$ between their rendered images $\mathcal{R}(\hat{G}^{t}),\mathcal{R}(G^{t})$ respectively:
$\min_{\mathcal{F}_{\text{deform}}}\mathcal{L}_{\text{CD}}(\hat{G}^{t},G^{t})+\mathcal{L}_{\text{photo}}(\mathcal{R}(\hat{G}^{t}),\mathcal{R}(G^{t}))$ (3)
where $\hat{G}^{t}=\{\hat{\mathbf{g}}_{i}^{t}\}_{i=1}^{N^{s}}$ represents the deformed source Gaussians, with each point $\hat{\mathbf{g}}^{t}$ obtained by applying the transformation predicted by $\pi=\mathcal{F}_{\text{deform}}(\bm{\mu})$ to the corresponding source point $\mathbf{g}^{s}\in G^{s}$, where $\pi$ is a similarity transformation containing rotation $\mathbf{R}$, translation $\mathbf{t}$, and scaling $s$. Specifically, to obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, we first compute the transformation for each point $\mathbf{g}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$ using Eq. 1, and apply the transformation to the means and rotation of the source Gaussian points. With $\hat{\bm{\mu}}^{t}=s\mathbf{R}\bm{\mu}^{s}+\mathbf{t}$ and $\hat{\mathbf{q}}^{t}=\mathbf{q}^{s}\otimes\mathbf{q}(\mathbf{R})$, we can obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, where $\otimes$ denotes quaternion multiplication and $\mathbf{q}(\mathbf{R})$ converts the rotation matrix $\mathbf{R}$ to its quaternion representation. For simplicity, we denote this process as $\hat{\mathbf{g}}=\pi\circ\mathbf{g}$.
Pipeline Corr. Foldchair Fridge Laptop Oven Scissor Stapler USB Washer Blade Storage
Motion Ang Err $\downarrow$ AGS[16] - 62.083 36.568 86.058 78.129 62.083 54.613 88.677 87.454 59.562 75.744
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 0.974 13.267 failed failed 5.209 failed 12.603 35.926 3.716 49.799
Ours Ours 0.095 0.914 1.020 2.365 0.839 1.337 1.872 1.737 0.878 0.525
Pos Err $\downarrow$ AGS[16] - 0.032 0.124 1.805 0.252 0.032 0.149 1.209 0.231 - -
Ours GaussReg[3] failed failed failed failed failed failed failed failed - -
Ours FM-Dense[2] 0.729 0.658 failed failed 0.394 failed 0.281 0.417 - -
Ours Ours 0.180 0.007 0.106 0.042 0.058 0.196 0.099 0.157 - -
Motion Dist $\downarrow$ AGS[16] - 70.900 75.548 46.925 29.381 70.900 64.865 61.955 33.345 1.275 0.213
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 13.590 25.261 failed failed 3.652 failed 11.285 17.509 0.089 0.192
Ours Ours 0.861 2.279 0.788 2.654 1.516 0.930 1.877 2.926 0.017 0.118
Rendering PSNR $\uparrow$ AGS[16] - 9.349 11.294 13.787 13.871 13.481 12.898 15.856 17.644 24.543 14.333
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 17.897 19.256 21.161 19.356 23.276 19.002 22.517 20.632 25.741 20.900
Ours Ours 24.659 26.920 26.491 23.434 28.823 26.532 27.016 27.416 29.041 25.493
SSIM $\uparrow$ AGS[16] - 0.652 0.722 0.809 0.723 0.846 0.831 0.842 0.877 0.962 0.765
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 0.889 0.916 0.923 0.893 0.953 0.930 0.821 0.933 0.970 0.896
Ours Ours 0.940 0.963 0.955 0.931 0.967 0.953 0.956 0.966 0.974 0.927
Geometry CD $\downarrow$ AGS[16] - 1188.824 717.780 893.149 244.862 1188.824 1599.261 467.633 268.377 90.288 231.355
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 4.814 6.399 failed failed 0.501 failed 8.935 25.515 2.789 15.599
Ours Ours 0.856 2.051 0.610 6.449 0.264 1.216 1.430 1.615 1.788 6.443
Table 1: Comparisons to the baseline methods with 4-view images from each articulation states (source and target). Objects with $\dagger$ are with prismatic joints.
3.3 Joint Optimization
To obtain an initial segmentation and transformations, we leverage TEASER++ [59], a robust pose estimation algorithm. For clarity, we first describe the common two-part (root/leaf) scenario (see Sec. 3.4 for multi-part). We apply TEASER++ to the correspondences from Sec 3.2. It finds the largest rigid component (inliers) as the root $r$, treating the remaining outliers as the leaf $l$. This yields initial transformations $\pi_{r},\pi_{l}$ and part labels $p^{s}$.
Both the initial geometry from FreeSplatter and the part-pose estimates from TEASER++ are coarse approximations. Therefore, a joint optimization is essential to mutually refine the geometry, segmentation, and articulation parameters. To do this without the optimization getting stuck in a bad local minimum, we introduce a three-phase staged refinement strategy. This coarse-to-fine approach first stabilizes the large-scale rigid motions (Phase 1), then refines the underlying geometry and segmentation (Phase 2), and finally polishes the visual appearance (Phase 3).
We define our full learnable model and its parameters. We initialize the learnable Gaussian splats $\bar{G}^{s}=\{\bar{\mathbf{g}}^{s}_{i}\}_{i=1}^{N^{s}}$, where each $\bar{\mathbf{g}}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s},p^{s})$ is augmented with its initial part label $p^{s}$. We also treat the rigid transformations $\pi_{r}$ and $\pi_{l}$ as learnable parameters. To compensate for non-rigidity and initial estimation errors, we introduce a learnable, per-Gaussian displacement term $\delta\bm{\mu}$. To simplify the notation, we define an auxiliary function $f(\bar{\mathbf{g}}^{s}_{i},\delta\bm{\mu}_{i},\pi)$. This function first selects the appropriate rigid transformation $\pi_{p^{s}}$ from the set $\pi$ (using the part label $p^{s}$ contained within $\bar{\mathbf{g}}^{s}_{i}$), applies it, and then adds the displacement $\delta\bm{\mu}_{i}$ to the transformed mean. This yields the final deformed Gaussian splats $\bar{G}^{t}=\{\bar{\mathbf{g}}^{t}_{i}|\bar{\mathbf{g}}^{t}_{i}=f(\bar{\mathbf{g}}^{s}_{i},\delta\bm{\mu}_{i},\pi),\bar{\mathbf{g}}^{s}_{i}\in\bar{G}^{s}\}$.
Phase 1: Rigid Pose Refinement The goal of this phase is to stabilize the coarse poses from TEASER++. We apply two constraints: the source Gaussian geometry $\bar{G}^{s}$ is frozen, and the non-rigid displacement $\delta\bm{\mu}$ is set to zero. Using the fixed initial segmentation $p^{s}$, we only optimize the rigid transformations $\pi=\{\pi_{r},\pi_{l}\}$ by minimizing a joint loss:
$\mathcal{L}_{\text{joint}}=$ $\mathcal{L}_{\text{CD}}(\bar{G}^{t},G^{t})+$ (4)
$\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{t}),\mathcal{R}(G^{t}))+\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{s}),\mathcal{R}(G^{s}))$
The optimization objective for this phase is:
$(\pi_{r}^{*},\pi_{l}^{*})=\arg\min_{\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}$ (5)
Phase 2: Geometry and Occlusion Refinement With the rigid poses stabilized in Phase 1, we now address the inaccuracies in the underlying geometry and the initial part segmentation. In this phase, we unfreeze the parameters in $\bar{G}^{s}$ and activate the learnable displacement term $\delta\bm{\mu}$. This per-Gaussian term aims at compensating for any non-rigid errors, which typically caused by incorrect part segmentations. We encourage the displacement to be sparse by penalizing it with $\mathcal{L}_{\text{rigid}}=\sum_{i=1}^{N}\left\|\delta\bm{\mu}_{i}\right\|_{1}$.
Crucially, this displacement term also serves as a heuristic to correct mis-segmented points. As suggested in [10], we periodically update segmentation labels $p^{s}$. We assume points with a large $\delta\bm{\mu}$ are mis-assigned, and we re-assign their label $p$ based on the smallest rigid alignment error:
$p=\arg\min_{p\in\{r,l\}}\|\bar{\bm{\mu}}^{p}-\bar{\bm{\mu}}^{t}\|_{2}$ (6)
with $\bar{\bm{\mu}}^{p}$ for the means of $\bar{\mathbf{g}}^{p}=\pi_{p}\circ\bar{\mathbf{g}}^{s},p\in\{r,l\}$.
The full optimization objective minimizes the joint loss plus the rigidity penalty, updating all parameters (geometry, transformations, and displacements) simultaneously:
$\min_{\{\delta\bm{\mu}\},\bar{G}^{s},\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}+\mathcal{L}_{\text{rigid}}$ (7)
Phase 3: Pose and Appearance Refinement The previous phases refined the geometry and produced an as-rigid-as-possible part segmentation. In this final phase, we lock this geometry (freezing $\bar{G}^{s}$ and setting $\delta\bm{\mu}=\mathbf{0}$) to focus exclusively on finetuning the per-part pose estimation ($\pi_{r},\pi_{l}$) and visual quality (colors $\mathbf{c}$). This is supervised purely by the photometric loss against the original input images:
$\min_{\{\pi_{r},\pi_{l},\mathbf{c}\}}\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{s}),I^{s})+\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{t}),I^{t}).$ (8)
In practice, we use 500 iterations for the first phase, 2500 iterations for the second phase, and 2000 iterations for the final phase. Finally, we decompose the articulation parameters into axis $\mathbf{a}\in\mathbb{R}^{3}$, angle $\theta\in\mathbb{R}$, and pivot point $\mathbf{p}\in\mathbb{R}^{3}$ for revolute joints, while only axis $\mathbf{a}$ and magnitude $d\in\mathbb{R}$ for prismatic joints.
3.4 Modeling Multiple Moving Parts
Figure 3: Illustration of the extension to multiple parts. We progressively estimate the segmentation and articulation one part at a time with extra images for new target articulation states.
For simplicity, we have described our method for a two-part object. Our framework extends to objects with multiple parts by applying it iteratively, as shown in Fig. 3. This requires a common source image set ($\mathcal{I}^{s}$) and additional target sets ($\mathcal{I}^{t1},\mathcal{I}^{t2},\dots$), one for each moving part. We first run the full pipeline on the $(\mathcal{I}^{s},\mathcal{I}^{t1})$ pair to model the first part. We then repeat the process on the $(\mathcal{I}^{s},\mathcal{I}^{t2})$ pair to model the second part, and so on. The key assumption is that each source-target pair isolates the motion of only one part.
4 Experiments
Corr. Oven Storage USB Laptop
GaussReg Refer to caption Refer to caption Refer to caption Refer to caption
FM-Dense Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative analysis of correspondence methods. We visualize the part segmentations computed by our TEASER [59] solver, which is fed correspondences from each method. Inliers for the two main segmented parts are colored blue and orange; outliers and non-correspondences are shown in gray. Better view in color and zoom in.
4.1 Experimental Setup
Dataset.
We evaluate our method on the synthetic 3D PartNet-Mobility dataset [39], following the evaluation protocol established by previous work [30, 10, 16]. Our evaluation metrics encompass three key aspects: (1) For rendering, we use PSNR (Peak Signal-to-Noise Ratio) and SSIM (Structural Similarity Index) to evaluate the quality of the rendered images for unseen views in both source and target articulation states. (2) For articulation parameter estimation, we use the angular error, position error of the estimated motion axis and additional motion distance to evaluate the accuracy of articulation parameters as in previous work [10, 30, 16]. More details are provided in the supplementary material. (3) For 3D reconstruction precision, we use the Chamfer Distance (CD) to evaluate the precision of the estimated 3D geometry. For PSNR, SSIM and CD, we report the average values of the two articulation states.
GT AGS AGS-Full Ours
FM-Dense Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption failed Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption failed Refer to caption
Figure 5: Qualitative evaluation for novel view synthesis in target state. We can see from the results that both AGS-GT and AGS-VGGT fail to reconstruct the object in the setting of 4-view images. In the meanwhile, our method demonstrates similar rendering quality compared to AGS-Full, which is trained with 100 images per articulation state with ground truth camera poses.
Baselines.
To our knowledge, no existing work addresses self-supervised articulated object modeling from sparse, unposed images. Our evaluation is therefore twofold: a component-wise analysis to validate our correspondence method and a full-pipeline comparison against state-of-the-art methods. For the component analysis, we fix our pipeline and swap our correspondence module with two alternatives: (1) GaussReg [3], a pre-trained rigid GS registration method that provides 1024 sparse keypoint correspondences, and (2) FM-Dense [2], a pre-trained non-rigid matching network, to which we feed 20480 downsampled points from our GS to obtain dense correspondences. For the full-pipeline comparison, we evaluate against ArticulatedGS (AGS) [16] with the same number of images ($K=4$) and ground truth, aligned camera parameters. Besides, we use the VGGT [53] reconstructed point clouds to initialize the AGS for better reconstruction. We will make our code and models publicly available.
4.2 Comparison with Baselines
Quantitative Evaluation
Quantitative results are presented in Tab. 1.
We first compare our method against ArticulatedGS (AGS) [16]. The results show that AGS, even with GT aligned camera poses and VGGT initialization, fails to converge in a sparse 4-view setting. This is because its optimization, designed for dense data, introduces significant noise from sparse inputs, and the limited views are insufficient for its motion estimation to converge. These factors lead to catastrophic errors, demonstrating that existing SOTA methods are not robust to this sparse-input scenario.
Our qualitative analysis in Fig. 4 visually confirms our quantitative findings by showing the output of the TEASER [59] solver for each baseline (inliers in blue/orange, outliers in gray). The GaussReg [3] output is entirely gray, confirming its 1024 sparse, noisy correspondences are insufficient to initialize segmentation. FM-Dense [2] proves highly unstable; the visualization shows most of its 20k correspondences are rejected as outliers, and even when a solution is found (e.g., Laptop), segmentation is incorrect, with inliers (blue/orange) assigned to the same part. In contrast, our method produces a dense, reliable map with a high percentage of correctly segmented inliers. A small volume of outliers is visible at part boundaries (e.g., USB, Storage), which are refined in the joint optimization step.
RGB input AGS-GT AGS-Full FM-Dense Ours GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Visualization of 3D Gaussian points with part-level segmentation and articulation axes. Ground truth axises are shown in green, predictions in red. AGS-Full denotes the AGS pipeline with dense and posed images input. Better view in color and zoom in.
$k_{0}=10^{-5}$ $k_{0}=10^{-8}$ $k_{0}=10^{-11}$
Refer to caption Refer to caption Refer to caption
Coarse segmentation at source state
Refer to caption Refer to caption Refer to caption
Deformation Overlay of Gaussian Points
Figure 7: Ablation study for the choice of $k_{0}$ for the deformation field. For segmentation, different color denotes different parts identified using TEASER at source state. For deformation overlay, red points denote the deformed Gaussian points $\hat{G}^{t}$, and green points denote the Gaussian points at target state $\hat{G}^{t}$.
RGB input Novel articulation synthesis Part-level seg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Real-world objects. We show the input images (left), novel articulation synthesis results (middle), and part-level segmentation with articulation axes (right). The first three objects are single-part articulated objects, while the last one is a multi-part articulated object. For the multi-part object, we collect three sets of input images per articulation state. The segmentation results are shown in different colors for each part. The green lines denote the estimated articulation axes.
Qualitative Evaluation
We provide a qualitative analysis of the correspondence methods in Fig. 4. This figure visualizes the output of the TEASER [59] solver when fed correspondences from each baseline. Inliers for the two main parts are colored blue and orange, while outliers and non-correspondences are gray. The results visually confirm our quantitative findings. The GaussReg [3] baseline fails completely; its point cloud is rendered entirely gray, confirming that its 1024 sparse and noisy keypoint correspondences are insufficient to initialize segmentation. FM-Dense [2] proves highly unstable; the visualization shows most of its 20k correspondences are rejected as outliers, and even when a solution is found (e.g., Laptop), segmentation is incorrect, with inliers (blue/orange) assigned to the same part. In contrast, our method produces a dense, reliable map with a high percentage of correctly segmented inliers onto the distinct rigid parts. A small volume of outliers (gray points) is visible at the part boundaries on objects like the USB and Storage, which will can be further optimized by the joint optimization step.
Further qualitative evidence is provided in Fig. 6, which visualizes the final part segmentation and estimated joint axes. For this comparison, we include AGS [16] (using its full, dense-view and posed-image configuration) as a state-of-the-art upper bound. As the figure illustrates, our method’s joint parameter estimation is visually on par with the SOTA AGS baseline, despite our significantly more challenging sparse and unposed input. In contrast, the FM-Dense [2] baseline demonstrates degraded performance, resulting in clear errors in both segmentation and joint estimation. This visualization further highlights the robustness and accuracy of our full pipeline. We also provide the qualitative evaluation for novel articulation synthesis in Fig. 5. For more visualizations, please refer to the supplementary.
4.3 Ablation Studies
Effect of $k_{0}$ for deformation field
We find that our method is generally robust to the choice of $k_{0}$ in the deformation field. To analyze its effect, we conduct experiments with different $k_{0}$ values, with qualitative results shown in Fig. 7. A smaller $k_{0}$ enforces near-rigid transformations across the whole object, leading to under-deformation and failure to reach the target state (e.g., right side of the second row). In contrast, a larger $k_{0}$ relaxes rigidity too much, causing over-deformation. Both extremes result in poor initial part segmentation and degraded performance. Empirically, we find that setting $k_{0}$ between $10^{-7}$ and $10^{-9}$ achieves a good balance, preserving rigidity while allowing sufficient flexibility for deformation.
Real-world Experiments
To demonstrate the generalization of our approach, we conduct experiments on four real-world objects (a bottle, projector, drawer, and lego toy), capturing four input images per articulation state. We employ the Segment Anything Model (SAM) [43] to remove backgrounds prior to processing. For multi-part objects, like the lego toy, we capture three image sets (one per state) as detailed in Sec. 3.4. The results in Fig. 8 show that our method successfully generates seamless reconstructions for these real-world objects, validating its practical applicability.
5 Conclusion
We have introduced PAOLI, the first framework to successfully learn high-fidelity articulated object representations from sparse, unposed images. We identified that the central challenge in this setting is not just reconstruction, but the 3D-to-3D correspondence problem between unaligned 3D models.
Our solution is a test-time, optimization-based deformation field that finds dense and robust correspondences by leveraging both 3D geometry and 2D rendering losses. Our experiments show this approach is highly effective, whereas prominent baselines—including sparse keypoint methods and pre-trained dense networks—fail to solve the problem. Our full pipeline successfully recovers high-fidelity geometry, appearance, and kinematics, demonstrating a robust path forward for articulated object learning in practical, unposed scenarios.
While effective, our method assumes rigid parts and requires at least two distinct articulation states with adequate view coverage. Furthermore, our alignment may fail if the initial, arbitrary coordinate systems of the 3D GS reconstructions are extremely misaligned. More detailed discussion on limitations can be found in the supplementary.
[…]
Supplementary Material
6 Supplementary Material
6.1 Implementation Details
Deformation Network.
The deformation network $\mathcal{F}_{\text{deform}}$, introduced in Sec. 3.2, is a 3-layer MLP with 128-dimensional hidden layers and ReLU activations. The network outputs a rigid transformation where rotation is parameterized as Euler angles.
Computation Efficiency
Our full pipeline runs in approximately 15 minutes per object on a system with an AMD Zen3 7950X CPU and an NVIDIA RTX 4090 GPU. The runtime breaks down as follows:
  • Step 1: $\sim$1 minute for initial reconstruction and camera pose refinement.
  • Step 2: $\sim$4 minutes for deformation field optimization.
  • Step 3: $\sim$10 minutes, which includes a few seconds for the TEASER++ solver and the remainder for joint refinement.
The peak memory usage is approximately 15 GB of VRAM during the Step 3.
6.2 Ablation Studies
Full Pipeline w.o. Cam. Opt. w.o. $L_{CD}$ w.o. $L_{photo}$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Visualization of the results after TEASER initialization in Step 3. The main impact of these ablated components occurs before this initialization step. Better view in color and zoom in.
Impact of Key Pipeline Components.
Table 2: Ablation study on the key components of our pipeline. We report the average metrics across all objects. Removing any component results in performance degradation, highlighting their importance.
Components Metrics
cam_opt $\mathcal{L}_{\text{CD}}$ $\mathcal{L}_{\text{photo}}$ Ang Err $\downarrow$ Pos Err $\downarrow$ Motion Dist $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$
$\checkmark$ $\checkmark$ $\checkmark$ 1.160 0.105 1.318 26.573 0.953 2.272
[…]
We conducted an ablation study to analyze the impact of key components in our pipeline: camera pose refinement in step 1 and the losses used for deformation network optimization ($\mathcal{L}_{\text{CD}}$ and $\mathcal{L}_{\text{photo}}$) in step 2. As shown in Tab. 2, removing any of these components leads to a degradation in performance.
These components are crucial for initializing the part segmentation and pose estimation in Step 3. To understand the performance drop, we examined the initialization results. As illustrated in Fig. 9, omitting these components leads to poorer initializations, which subsequently hampers the final joint optimization stage. This demonstrates that each component plays a vital role in the robustness of our framework.
Effectiveness of the Joint Optimization
The effectiveness of our joint optimization process is demonstrated in Fig. 10. The figure shows how the process corrects initial errors: despite dense matching, the state after TEASER initialization still has spatial discrepancies between parts. Phase 1 recovers connectivity, Phase 2 refines part segmentation, and Phase 3 produces the final, rigid, joint-optimized output.
TEASER init. Phase 1 Phase 2 Phase 3
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Visualization of the effectiveness of the joint optimization process, showing the progression from the initial state to the final refined output.
Effect of the number of input images.
With just 4-view images, traditional articulated object modelling methods struggle to reconstruct objects even when in static state. We investigate how the number of input views affects performance by varying the views per articulation state. Fig. 11 shows the rendering quality for ArticulatedGS [16] methods improves with more input images, saturating after 16 images (VGGT runs out of memory with 64 images on our 48GB VRAM GPU), even with VGGT-estimated poses. Our method matches the saturated performance without ground truth poses with only 4 input images and stays consistent across different view counts with qualitative evaluation.
More Qualitative Results
We provide additional qualitative results for correspondences and segmentation in Fig. 13 and Fig. 15, respectively. Besides, correspondences visualization on the real world images are also shown in Fig. 14.
6.3 Failure Cases
Extreme Initial Misalignment.
We demonstrate some failure cases in Fig. 12. Our method relies on FreeSplatter [62] for initial 3D reconstruction, which establishes a coordinate system based on the first input image of a set. If the first images from the source and target sets are captured from drastically different viewpoints (e.g., front vs. back), the resulting 3D models can be severely misaligned. This extreme misalignment can cause our deformation network to converge to a poor local minimum, especially for objects with geometric symmetries. For instance, as shown with the real-world storage drawer in Fig. 7, providing opposing side views as initial frames causes the network to favor a non-rigid deformation of the drawer rather than finding the correct global alignment. To mitigate this, we recommend selecting initial views from similar perspectives for each image set.
4-view 8-view 16-view
AGS-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AGS-VGGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Qualitative evaluation for the effect of the number of input images. We show novel view synthesis for source state (left) and target state (right) at each settings. From top to bottom: AGS-GT, AGS-VGGT, Ours.
Input reference view Render deformation target states 3D overlay
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Visualization of the failure cases in extreme misalignment case (top) and thin structures with large motion (bottom). The 3D overlay shows the itermediate state during the optimization of deformation field, the red is the deformed source state, the green is the target state.
Thin Structures with Large Motion.
The framework can struggle with objects that have very thin parts undergoing large articulation. A thin structure is represented by a smaller volume of Gaussian splats, which provides weaker rigidity constraints during the deformation step. When combined with a large motion, the deformation field may fail to preserve the part’s rigidity, leading to inaccurate correspondence and subsequent errors in joint estimation. An example of this can be seen in Fig. 12.
6.4 Evaluation Metrics
As we mentioned in the main paper, we will decompose the relative motion between leaf part and root part into axis $\bm{a}$, pivot $\bm{p}$ and angle $\alpha$ for revolute joints and axis $\bm{a}$ and magnitude $d$ for prismatic joints. Since our method is a pose-free method, we need to recover the reconstructed geometry in the ground truth coordinate system. The coordinate transformation can be computed by aligning the position of the estimated camera pose and the ground truth camera pose using Umeyama’s algorithm [52]. To further mitigate the effects of noisy pose estimation, we optimize the transformation matrix initialized with Umeyama’s algorithm by minimizing the chamfer distance between the reconstructed and ground truth geometries. We apply identical preprocessing procedures when evaluating pose-free baseline methods.
Ang Err:
The angular error computes the orientation difference between the predicted axis direction and the ground truth.
Pos Err:
The position error computes the minimum distance between the rotation axis and the ground truth which takes the position of the pivot point into account. Note that this error only applies to revolute joints.
Motion Dist:
For motion distance computes the difference in movements between estimated motion and ground truth. For revolute joints, we compute the geodesic distance between the predicted rotation matrix and the ground truth rotation matrix. For prismatic joints, we compute the distance between the predicted translation magnitude and the ground truth translation magnitude.
Ours FM-Dense [2] GaussReg [3] Ours FM-Dense [2] GaussReg [3]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Visualization of correspondences for our methods and baselines. Inliers for the two main rigid parts (from TEASER solver) are colored green and orange. For GaussReg [3], outliers are shown in black. Since our method and FM-Dense [2] produce dense correspondences, we visualize a sample of up to 2000 inliers per part. Better view in color and zoom in.
Source Input Target Input Correspondences
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 14: Visualization of correspondences for our methods and baselines on real-world data. Better view in color and zoom in.
Source/Tgt Novel View/Articulation Synthesis Seg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Additional qualitative results for novel view and articulation synthesis. For each object, we show one source and one target input view, followed by a sequence of renderings from our model with novel interpolated unseen articulation states. We also show the segmentation with estiamted asxis at the end of each row.